how many ounces of a 30% solution must be added to 50 ounces of a 12% saline solution to make a 20% saline solution? Show work please.
Answer by kittycat1613(12)

Let x= # of ounces of 30% solution
x+50= # of ounces of 20% saline solution
.3(x)+.12(50)=.2(x+50)
.3(x)+6=.2x+10
-.2x -6 -.2x -6
[.1x=4]*10
1x=40
x=40
Check
40(.3)+50(.12)=90(.2)
12+6=18
18=18
Correct
Answer: 40 ounces of a 30% solution must be added